### v1

# Dynamic modes of active Potts models with factorizable numbers of states

Abstract
We studied the long-term nonequilibrium dynamics of $q$-state Potts models with $q=4$, $5$, $6$, and $8$ using Monte Carlo simulations on a two-dimensional square lattice. When the contact energies between the nearest neighbors for the standard Potts models are used, cyclic changes in the $q$ homogeneous phases and $q$-state coexisting wave mode appear at low and high flipping energies, respectively, for all values of $q$. However, for a factorizable $q$ value, dynamic modes with skipping states emerge, depending on the contact energies. For $q=6$, a spiral wave mode with three domain types (one state dominant or two states mixed) and cyclic changes in three homogeneous phases are found. Although three states can coexist spatially under thermal equilibrium, the scaling exponents of the transitions to the wave modes are modified from the equilibrium values.
I Introduction
Various types of pattern formation have been observed in nonequilibrium systems [1, 2, 3, 4, 5, 6, 7]. Classically, they have been called self-organizing systems [1, 2], dissipative structures [1], and synagetics [2, 3], and recently, nonreciprocal systems [8, 9, 10, 11], featuring different aspects. In biological systems, spatiotemporal patterns have been observed from microscopic to macroscopic scales (cells and tissues to animal populations) [3, 7, 12, 13, 14, 15]. Under far-from-equilibrium conditions, the thermal fluctuations occurring in macroscopic patterns are typically negligible and the dynamics can be explained using deterministic continuum equations. In contrast, the thermal fluctuations can be significant in microscopic patterns, such as intracellular waves. However, the effects of the fluctuations and nucleation are not understood so far.
Lattice models, including Ising and Potts models, have been widely used to study the phase transitions occurring at thermal equilibrium [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The nonequilibrium dynamics have been investigated by imposing periodic external fields [31, 32] and spatial temperature inhomogeneity [33]. Recently, the cyclic flipping energies of three or four states have been introduced to these lattice models, and their pattern formations have been studied [34, 35, 36, 37, 38, 39]. We have used the standard Potts models [16, 17] for three [34, 35] and four [36] states in a two-dimensional (2D) square lattice under both cyclic symmetry [34, 36] and asymmetry [35, 36]; Hence, the transitions between dynamic modes, such as the cyclic changes of homogeneous phases and spiral waves, were revealed. These dynamics were also obtained in an off-lattice model for undulating fluid membranes [40]. These models exhibit stable spatiotemporal patterns in the long-term limit, unlike lattice Lotka–Volterra models, which exhibit an absorbing transition to a uniform state [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Manacorda and Fodor reported wave and global oscillation modes using a lattice model, which allows the occupancy of multiple particles at each site [37]. Further, Avni et al. have added a nonreciprocal interaction to the Ashkin-Teller model with a pairwise spin interaction (it is also called four-state vector-Potts and clock models [16]) and revealed changes in the scaling exponents of the second-order transition between disordered and ordered phases [38, 39]. However, this ordered phase maintains only for a short period (tproportional-$t\propto$ system size $N$) starting from a uniform initial state.
With the aim of clarifying the effects of thermal fluctuations on pattern formation and nonequilibrium phase transitions, this study investigates several transitions between the dynamic modes in the long-term limit ($t\to\infty$), which are independent of the initial states. We use general $q$-state Potts models with $q=4$, $5$, $6$, and $8$ under cyclically symmetric conditions. Various dynamic patterns are formed, particularly for a factorizable $q$ number. Although phases comprising factorized numbers of states have been reported at thermal equilibrium [21, 22, 23, 24, 55], to the best of our knowledge, spatiotemporal patterns with factorized symmetry have not been reported. We reveal several types of patterns with factorized symmetries and transitions between them.
The model and methods are described in Sec. II. Simulation results are presented and discussed in Sec. III. The results for $q=6$, $5$, $4$, and $8$ are described in Secs. III.1,  III.2III.3, and III.4, respectively. The relationship with other models is discussed in Sec. IV. Finally, a summary is presented in Sec. V.
II Active Potts Models
A $q$-state Potts model of a 2D square lattice with side length $L$ is considered. The total number of sites is $N=L^{2}$, and each site has a state $s\in[0,q-1]$. The nearest neighboring sites ($i$ and $j$) have contact energies $J_{s_{i},s_{j}}$:
$H_{\mathrm{int}}=-\sum_{\langle ij\rangle}J_{s_{i}s_{j}}.$ (1)
In standard Potts models, $J_{s_{i},s_{j}}=J_{0}\delta_{s_{i},s_{j}}$ is used. In equilibrium systems, each state can additionally have a self-energy $\varepsilon_{s}$, and the forward and backward flip rates is $\exp(-\Delta H_{s_{i}s^{\prime}_{i}})$ for flipping of a single site from $s$ to $s^{\prime}$. The thermal energy $k_{\mathrm{B}}T$ is normalized to unity, $\Delta H_{s_{i}s^{\prime}_{i}}=\Delta H_{\mathrm{int}}-h_{s,s^{\prime}}$, and the flipping energy $h_{s,s^{\prime}}=\varepsilon_{s}-\varepsilon_{s}^{\prime}$. The cyclic sum of the flipping energy is $\sum_{k=0}^{q}h_{k,[k+1]}=0$ at equilibrium, where $[k]$ indicates $k\bmod q$. We extended this model to a nonequilibrium situation, in which $\sum_{k=0}^{q}h_{k,[k+1]}>0$, but $h_{ss^{\prime}}=-h_{s^{\prime}s}$ remains [34]. Hence, the detailed balance can be locally satisfied for flips between $s$ and $s^{\prime}$, but not globally for cycles ($s=0\to 1\ ...\to q-1\to 0$). For $q=3$, this corresponds to the rock–paper–scissors relationship. Note that this situation can be realized by reactions on a catalytic surface [56, 57, 58, 59, 60] and molecular transport through a membrane [61, 62, 63, 64, 40]. In the context of surface reactions, the state $s=0$ represents an unoccupied site, $s=1$ is a reactant-bound state, and $s=2$ to $q-2$ and $s=q-1$ are intermediate and final products, respectively. The bulk reaction energy for the final product is expressed as $\sum_{k=0}^{q}h_{k,[k+1]}$. In the case of molecular transport, the molecules have $q-1$ intermediate states in the membrane.
Figure 1: Six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$ and $L=128$. (a–d) Snapshots. The yellow, green, light-blue, dark-blue, magenta, and red sites (light to dark in grayscale) represent $s=0$, $1$, $2$, $3$, $4$, and $5$, respectively. (a) Homogeneous cycling of six states (HC6) at $h=0.8$ and $J_{k,[k+2]}=-1$. (b) Waves of six states (W6) at $h=1.5$ and $J_{k,[k+2]}=-1$. (c) Waves of three states (W3) at $h=1$ and $J_{k,[k+2]}=0.5$. (d) Mixing of six states (M6) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, and magenta diamonds represent HC6, W3, and M6, respectively. The blue up-pointing triangles and green down-pointing triangles represent W6 and the intermediate wave mode (WI), respectively.
Figure 2: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}=-1$ and $L=128$. (a) W6 at $h=1.5$. (b) WI at $h=1.2$. (c) HC6 at $h=0.8$. (c, inset) Enlarged density development at $t\sim 370~{}000$ showing skipping of $s=3$ and $4$ phases.
Figure 3: Dynamics of the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$. (a–b) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states as functions of $h$ at (a) $J_{k,[k+2]}=-1$ and (b) $J_{k,[k+2]}=0$. The solid and dashed lines represent the data at $L=128$ and $256$, respectively. The bidirectional arrows at the top represent the ranges of dynamic modes at $J_{k,[k+2]}=-1$. (c) System-size dependence of phase boundaries. The solid and dashed lines represent the data at $J_{k,[k+2]}=-1$ and 00, respectively.
Figure 4: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$. (a) W3 at $J_{k,[k+2]}=0.5$. (b) Transition point at $J_{k,[k+2]}=1.214$. (c) M6 at $J_{k,[k+2]}=1.5$.
This study focuses on the dynamics occurring under the cyclic symmetry, such that the flipping energies are constant as $h_{s,s^{\prime}}=h$ and the contact energy $J_{s,s^{\prime}}$ depends on the state distance $|s-s^{\prime}|$ only. A site is randomly selected and the flip to the neighboring states ($s=k\to[k+1]$ or $[k-1]$) is performed using the Metropolis Monte Carlo (MC) algorithm. This flip is performed $N$ times per MC step (time unit). We have checked that the choice of the Metropolis and Glauber rates for the update process introduces no qualitative change in the dynamics at $q=3$ [34].
Through this study, $J_{k,k}=2$ and $J_{k,[k+1]}=0$, with $h$ and the other contact energies being varied: $J_{k,[k+2]}$ for $q=4$ and $5$, $J_{k,[k+2]}$ and $J_{k,[k+3]}$ for $q=6$, $J_{k,[k+2]}$, $J_{k,[k+3]}$, and $J_{k,[k+4]}$ for $q=8$. For the $q$-state standard Potts models in 2D under thermal equilibrium, the disorder-order transition is continuous and first-order for $q\leq 4$ and $q>4$, respectively, and the transition point is given by $J_{0,\mathrm{c}}=\ln(1+\sqrt{q})$ [16, 18]. Hence, the contact energy in this study, $J_{k,k}=2$, is sufficiently high for ordered phases in the standard Potts models to be obtained. The statistical errors were calculated from three or more independent runs (ten runs were used to calculate the scaling exponents).
III Results and Discussion
III.1 Six-State Potts Model
III.1.1 Modes for $J_{k,[k+3]}=J_{k,[k+2]}$
First, we describe the dynamics of the six-state Potts model ($q=6$) with $J_{k,[k+3]}=J_{k,[k+2]}$ (Figs. 16). Five modes are obtained as shown in Fig. 1.
[…]
Figure 5: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs.
Figure 6: Transition between W3 and M6 modes at $J_{k,[k+3]}=J_{k,[k+2]}$ for (a–c) $h=0$ and (d–f) $h=1$. Binder cumulant $U_{3}$, susceptibility $\chi_{3}$, and order parameter $R_{3}$ for three-fold rotational symmetry are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=1.725$ and $1.214$ for $h=0$ and $1$, respectively.
Figure 7: Six-state Potts model with $J_{k,[k+3]}=0$ and $L=128$. (a) Snapshot of homogeneous cycling of three states (HC3) at $h=1$ and $J_{k,[k+2]}=0.95$. (b) Snapshot of mixing of three states (M3) at $h=1$ and $J_{k,[k+2]}=1$. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC6, HC3, M3, W6, and WI, respectively.
Figure 8: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=0$ and $h=1$. (a) HC3 at $J_{k,[k+2]}=0.976$. (b) M3 at $J_{k,[k+2]}=1$.
[…]
For $J_{k,[k+2]}>0$, two additional modes emerge: wave modes of three states (W3) and a mixed phase of six states (M6), as shown in Figs. 1(c–e) and 4. As $J_{k,[k+2]}$ increases, the length of the boundary between the $s=k$ and $[k+2]$ domains increases owing to a reduction in the interfacial tension (see Fig. 5). When $J_{k,[k+2]}\simeq J_{k,k}=2$, the energy losses due to the contact between the different states (excluding the neighboring states $s=k$ and $[k+1]$) become negligibly small; hence, in the M6 mode, the six states are mixed to gain entropy. Note that this mode corresponds to the disordered phase at thermal equilibrium. For the intermediate $J_{k,[k+2]}$, three odd- or even-numbered states ($s=0,2,4$ or $s=1,3,5$) form domains exhibiting spiral waves, as observed in the three-state Potts model [34, 35] (see Fig. 1(c) and Movie S2). The other three states are present to small degrees at the domain boundaries. The boundary of the $s=k$ and $[k+2]$ domains moves via two-step flips ($s=k\to[k+1]\to[k+2]$), in which the clusters of $s=[k+1]$ states are swallowed up by the $[k+2]$ domain before growing. Waves of odd- and even-numbered states stochastically switch (see Fig. 4(a) and (b)). Note that this mode can be elucidated via symmetry factorization [55]. Static factorized phases for $q=2^{n}$ have been reported for the standard Ashkin-Teller model ($n=2$[21, 23, 24] and extended $n$-color models [22]. Here, the factorized (three-fold) symmetry remains within a $6=2\times 3$ symmetric system. For the wave propagation modes, the three-fold symmetry helps maintain their spiral centers [36].
Figure 9: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=0$ and $L=128$ for (a–b) $h=1$ and (c) $h=0.8$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs. (c) Lifetimes of one-phase state ($\tau_{1}$) and HC3 mode ($\tau_{\mathrm{HC3}}$). The bidirectional arrows at the top represent the ranges of dynamic modes.
The transition between the W3 and M6 modes is continuous. In the middle of the W3 range (Fig. 4(a)), the mean densities of the three dominant states are approximately $1/3$, and those of the other three states are approximately 00. With increasing $J_{k,[k+2]}$, the mean densities of the dominant and minor states decrease and increase, respectively. In the M6 mode, these mean densities converge to $1/6$ (see Fig. 4(b) and (c)). The transition between three and six states also occurs at thermal equilibrium ($h=0$), and the transition point $J_{k,[k+2]}=J_{\mathrm{c}}$ is weakly dependent on $h$ (see Fig. 1(e)). The transition is characterized by the order parameter $R_{3}$ for the three-fold symmetry. $R_{n}$ is defined as
$R_{n}$ $=$ $\langle s_{n}\rangle,$ (2)
$s_{n}$ $=$ $\frac{1}{N}\bigg{|}\sum_{j}^{N}\exp\Big{(}\frac{2n\pi{\mathrm{i}}s_{j}}{q}\Big{)}\bigg{|},$ (3)
where $n$ is a positive integer and delimited-⟨$\langle$..$\rangle$ … ⟩ is the long-term time average. When all the sites are occupied by one state, $R_{n}=1$ for any $n$. When the sites are equally occupied by the three odd-numbered states, $R_{1}=0$ and $R_{3}=1$. When the sites are equally occupied by all six states, $R_{1}=R_{3}=0$. The susceptibility $\chi_{n}$ and Binder cumulant $U_{n}$ are defined, respectively, as [20, 39]
$\chi_{n}$ $=$ $N(\langle{s_{n}}^{2}\rangle-\langle s_{n}\rangle^{2}),$ (4)
$U_{n}$ $=$ $1-\frac{\langle{s_{n}}^{4}\rangle}{3\langle{s_{n}}^{2}\rangle^{2}}.$ (5)
For second-order transitions, the scaling relations of the order parameter (here, $R_{3}$) are obtained in the vicinity of the transition point $J_{\mathrm{c}}$: $R_{3}L^{\beta/\nu}=\tilde{R}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}}]$, $\chi_{3}L^{-\gamma/\nu}=\tilde{\chi}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, and $U_{3}=\tilde{U}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, where $\tilde{R}_{3}(x)$, $\tilde{\chi}_{3}(x)$, and $\tilde{U}_{3}(x)$ are the scaling functions and $\nu$, $\beta$, and $\gamma$ are the critical exponents. Figure 6 shows the normalized $U_{3}$, $\chi_{3}$, and $R_{3}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We fitted each curve with a quartic function and calculated the scaling exponents by least-squares fits to these functions for $L\leq 90$ in the vicinity of the transition point. We obtained the following values: $J_{\mathrm{c}}=1.725\pm 0.002$, $1/\nu=0.95\pm 0.02$, $\gamma/\nu=1.742\pm 0.002$, and $\beta/\nu=0.1233\pm 0.0004$ for $h=0$; and $J_{\mathrm{c}}=1.214\pm 0.001$, $1/\nu=1.12\pm 0.02$, $\gamma/\nu=1.732\pm 0.007$, and $\beta/\nu=0.132\pm 0.002$ for $h=1$. The exponent $\nu$ is close to unity and is smaller for $h=1$ than for $h=0$ (equilibrium). Therefore, the exponents for the transition from M6 to the wave mode (W3) are modified from the equilibrium values between the mixed phases. Note that, in Fig. 6, for $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}\gtrsim 10$, the curves do not converge well either for $h=0$ or $1$. This deviation may be caused by the crossover effects [21] from other transitions (e.g., the W6–W3 transition).
Figure 10: Transition between HC3 and M3 modes at $J_{k,[k+3]}=J_{k,[k+2]}$ for (a–c) $h=0$ and (d–f) $h=1$. Binder cumulant $U_{1}$, susceptibility $\chi_{1}$, and order parameter $R_{1}$ on the single-phase occupancy are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=0.992$ and $0.977$ for $h=0$ and $1$, respectively.
III.1.2 Modes for $J_{k,[k+3]}=0$
Next, we investigate the $J_{k,[k+2]}$ dependence while keeping $J_{k,[k+3]}=0$ (Figs. 710). Although the dynamics for $J_{k,[k+2]}\leq 0$ are similar to those for $J_{k,[k+3]}=J_{k,[k+2]}$, two new modes (HC3 and M3) are found for $J_{k,[k+2]}>0$, instead of the W3 and M6 modes.
At $J_{k,[k+2]}=J_{k,k}=2$, three odd-numbered states ($s=1,3,5$) can be randomly mixed without lose of contact energy in the absence of the even-numbered states; similarly for the even-numbered states. By contrast, mixing of the six states is inhibited by the repulsion between the diagonal pairs ($s=k$ and $[k+3]$) because $J_{k,[k+3]}=0$. Thus, the mixed phases of the three states (M3) are stable for $J_{k,[k+2]}\simeq 2$, for which the boundaries of the clusters of each state are not smooth and moves diffusively rather than ballistically (see Figs. 7(b), (c), and 8(b), and Movie S3). No switching between the odd- and even-numbered states was observed during the available simulation periods $\sim 10^{9}$. This phase also exists at $h=0$.
For the intermediate values of $J_{k,[k+2]}$, homogeneous cycling of three phases ($s=1\to 3\to 5$ or $s=0\to 2\to 4$) occurs (HC3). The $s=k$ dominant phase involves small clusters of $s=[k+2]$ and $[k+4]$ states (see Figs. 7(a) and 8(a)). Hence, the $s=[k+1]$ dominant phase is skipped, because the $s=[k+2]$ phase grows rapidly from these clusters during the nucleation and growth of the $s=[k+1]$ phase.
The transition between the HC3 and M3 modes is continuous. As $J_{k,[k+2]}$ increases in the HC3 range, the densities of the dominant phase and the other two phases of the three major states decreases and increase, respectively (see Fig. 8). Additionally, the contact probability $p_{\mathrm{contact}}$ between these three states increases (see Fig. 9(b)). Hence, the three densities converge as the system approaches the M3 mode. This change can be characterized by the order parameter $R_{1}$. We calculated the scaling exponents using the method described in Sec. III.1.1. Figure 10 shows the normalized $U_{1}$, $\chi_{1}$, and $R_{1}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We obtained the following values: $J_{\mathrm{c}}=0.992\pm 0.001$, $1/\nu=1.24\pm 0.05$, $\gamma/\nu=1.75\pm 0.02$, and $\beta/\nu=0.137\pm 0.002$ for $h=0$; and $J_{\mathrm{c}}=0.977\pm 0.001$, $1/\nu=1.19\pm 0.04$, $\gamma/\nu=1.761\pm 0.008$, and $\beta/\nu=0.126\pm 0.005$ for $h=1$. These values are close to the $\nu=5/6$, $\gamma=13/9$, and $\beta=1/9$ obtained for the three-state standard Potts model [16]. The differences in the exponents for $h=0$ and $1$ are in the ranges of statistical errors, thus, the scaling exponents are not or little modified from the equilibrium values. This outcome may reflect the static nature of the two modes. That is, in most periods of the HC3 mode, a single-state dominant phase exists, similar to that obtained at equilibrium for $h=0$. The M3 mode is almost identical to the mixed phase at equilibrium. To obtain a distinct change in the scaling exponents, dynamic modes, such as wave propagation in W3, may be required.
Under low $h$, the HC3 mode changes to the HC6 mode with decreasing $J_{k,[k+2]}$. This change is continuous and can be characterized by the ratio of the lifetimes (one vs. three phases) and the distribution of the state densities. The state-density probability distribution $P(N_{s}/N)$ has three peaks in the HC3 mode (see Fig. 8(a)); in the HC6 mode, the two smaller peaks merge into a single peak. The lifetime $\tau_{1}$ of one phase is the average of the periods, in which one of the states ($s\in[0,q-1]$) maintains the condition of $N_{s}/N>0.95$. The HC3 mode of the odd numbers is taken to be initiated when $N_{\mathrm{odd}}/N$ exceeds $0.95$ and to end when $N_{\mathrm{even}}/N$ exceeds $0.95$, where $N_{\mathrm{odd}}=N_{1}+N_{3}+N_{5}$ and $N_{\mathrm{even}}=N_{0}+N_{2}+N_{4}$. Similarly, for that of the even numbers. The lifetime $\tau_{\mathrm{HC3}}$ of the HC3 mode is the average of these two periods. With increasing $J_{k,[k+2]}$, $\tau_{1}$ gradually increases. However, $\tau_{\mathrm{HC3}}$ rapidly increases for $J_{k,[k+2]}\gtrsim 0.3$ at $h=0.8$ (see Fig. 9(c)). We consider that the HC3 mode is determined by $\tau_{\mathrm{HC3}}/\tau_{1}>10$, since $\tau_{\mathrm{HC3}}$ is longer than $\tau_{1}$ in the HC3 mode.
Figure 11: Six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$. (a) Snapshot of spiral waves of three types of two-state mixed domains (M2W3) at $h=1.3$ and $J_{k,[k+3]}=2$. (b) Snapshot of homogeneous cycling of mixed phases (M2HC3) at $h=1.3$ and $J_{k,[k+3]}=2.5$. Two diagonal states ($s=2$ and $5$) are mixed. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, green down-pointing triangles, and light-blue squares represent HC6, M2W3, M2HC3, W6, WI, and W3, respectively.
Figure 12: Dependence on $J_{k,[k+3]}$ in the six-state Potts model at $J_{k,[k+2]}=0$ and $h=1.3$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs. (c) Binder cumulant $U_{3}$ for three-fold symmetry and cluster size ratio $d_{\mathrm{cl}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.
Figure 13: Time development of number densities of states in the six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$. (a) Temporal coexistence of M2W3 and M2HC3 at $h=1.3$ and $J_{k,[k+3]}=2.4$. (b) Temporal coexistence of W3 and HC6 at $h=0.7$ and $J_{k,[k+23]}=-0.6$.
III.1.3 Modes for $J_{k,[k+2]}=0$
Next, we investigate the $J_{k,[k+3]}$ dependence while keeping $J_{k,[k+2]}=0$ (Figs. 1113). For $J_{k,[k+3]}<0$, the W3 mode emerges in addition to the HC6, WI, and W6 modes, since the contact between the $k$ and $[k+2]$ states becomes more stable than that between the $k$ and $[k+3]$ states (see Figs. 11(c) and 12). The transition between the W3 and HC6 modes occurs via the temporal coexistence of the two modes, as shown in Fig. 13(b), similar to that occurring for the transition from HC3 to W3 in the three-state Potts model [34, 35]. The ratio of the W3 modes monotonously increases with decreasing $J_{k,[k+3]}$, and we determine the modes from the ratio of the phase periods.
[…]
In the M2W3 mode, the three types of mixed domains form spiral waves (see Fig. 11(a) and Movie S4). To characterize this mode, we calculated the $p_{\mathrm{contact}}$ values and cluster size ratio $d_{\mathrm{cl}}$. Clusters of single or diagonal two states are considered. Note that, in the latter case, when neighboring sites have the same state or diagonal state ($s_{i}=k$ and $s_{j}=k$ or $[k+3]$), they belong to the same cluster. With increasing $J_{k,[k+3]}$, the ratio $d_{\mathrm{cl}}=n_{\mathrm{c1}}/n_{\mathrm{c2}}$ of the mean sizes for these two clusters ($n_{\mathrm{c1}}$ and $n_{\mathrm{c2}}$ for the single and two diagonal states, respectively) decreases and $p_{\mathrm{contact}}$ of the diagonal pairs ($s=k$ and $[k+3]$) increases (see Fig. 12(b) and (c)). We consider that the W6 mode changes to the M2W3 mode when the the diagonal-pair $p_{\mathrm{contact}}$ exceeds the other probabilities.
Figure 14: Dependence on $J_{k,[k+3]}-J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+2]}+J_{k,[k+3]}=1$, $h=1$, and $L=128$. (a) Order parameters $R_{1}$ and $R_{3}$. The bidirectional arrows at the top represent the ranges of dynamic modes, including the coexistence states. (b) Susceptibilities $\chi_{1}$ and $\chi_{3}$. The conditions of $J_{k,[k+3]}-J_{k,[k+2]}=-1$, 00, and $1$ correspond to the points in the phase diagram of Figs. 7, 1, and 11, respectively. (c) Time development of number densities of states at $J_{k,[k+3]}-J_{k,[k+2]}=-0.4$. The HC3 and W3 modes temporally coexist.
Figure 15: Probabilities $p_{\mathrm{phase}}$ of one-phase dominant and multi-phase coexistence states as functions of $h$ at $J_{k,k^{\prime}}=2\delta_{k,k^{\prime}}$ and $L=128$ for the $q$-state Potts models with (a) $q=8$, (b) $q=5$, and (c) $q=4$. The plot in (c) is reproduced from Ref. 36. Licensed under CC BY.
III.1.4 Modes for $J_{k,[k+2]}+J_{k,[k+3]}=1$
Several modes are shown in the three phase diagrams (Figs. 1(e), 7(c), and 11(c)). At the final part of this analysis of the six-state Potts model, the intermediate conditions between these phase diagrams are considered. In detail, $J_{k,[k+3]}-J_{k,[k+2]}$ is varied while $J_{k,[k+2]}+J_{k,[k+3]}=1$ is maintained, as shown in Fig. 14. As a result, the M3 mode obtained at $J_{k,[k+3]}-J_{k,[k+2]}=-1$ ($J_{k,[k+2]}=1$ and $J_{k,[k+3]}=0$) sequentially changes to the HC3 mode at $J_{k,[k+3]}-J_{k,[k+2]}=-0.8$ ($J_{k,[k+2]}=0.9$ and $J_{k,[k+3]}=0.1$), the W3 mode at $J_{k,[k+3]}-J_{k,[k+2]}=0$ ($J_{k,[k+2]}=0.5$ and $J_{k,[k+3]}=0.5$), and the M2W3 mode at $J_{k,[k+3]}-J_{k,[k+2]}=1$ ($J_{k,[k+2]}=0$ and $J_{k,[k+3]}=1$). The transitions between M3 and HC3 and between HC3 and W3 are characterized by the order parameter $R_{1}$. The M3–HC3 transition is second-order, as described in Sec. III.1.2. At the HC3–W3 transition, the two modes temporally coexist (Fig. 14(c)), and the susceptibility $\chi_{1}$ has a broad peak (Fig. 14(b)), like in the first-order transition at equilibrium. The W3–M2W3 transition is likely second-order and is characterized by the order parameter $R_{3}$. The latter two transitions (HC3–W3 and W3–M2W3) have not seen the aforementioned phase diagrams. The transitions between other modes (e.g., M3–W3) may be revealed when the full parameter space is explored.
III.2 Five-State Potts Model
[…]
The dynamic phase diagram of the five-state Potts model is shown in Fig. 16. Since five is a prime number, no factorized-symmetry modes exist. At $J_{k,[k+2]}\simeq J_{k,k}=2$, a five-state mixed phase (M5) exists, as for the six-state Potts model. To characterize the transition between the M5 and W5 modes, we calculated $p_{\mathrm{phase}}$, $p_{\mathrm{contact}}$, and the mean cluster size $n_{\mathrm{c1}}$, as shown in Fig. 17; At $J_{k,[k+2]}\simeq 0.6$, a stepwise change is observed for $p_{\mathrm{phase}}$ of five states. Additionally, a minimum is observed for $p_{\mathrm{contact}}$ of $k$$[k+1]$ pairs and a sharp decrease in $n_{\mathrm{c1}}$ is initiated. For $J_{k,[k+2]}\lesssim 0.6$, $n_{\mathrm{c1}}$ is dependent on the system size; however, no size dependence is detected for $J_{k,[k+2]}\gtrsim 0.6$. In contrast, no size dependence is obtained for $p_{\mathrm{contact}}$ in either range (compare the curves for $L=128$ and $256$ in Fig. 17(b) and (c)). Thus, we consider $J_{k,[k+2]}\simeq 0.6$ is the transition point for $h=1$. The minimum of $p_{\mathrm{contact}}$ is used to determine the transition points for other values of $h$, since it is clearer than the other two features.
Figure 16: Dynamic phase diagram of the five-state Potts model at $L=128$. The red circles, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC5, M5, W5, and WI, respectively.
Figure 17: Dependence on $J_{k,[k+2]}$ in the five-state Potts model at $h=1$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$ and $k$$[k+2]$ pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.
III.3 Four-State Potts Model
We first describe phase behavior of the four-state Potts model at thermal equilibrium. The four-state Potts model corresponds to the standard (two-color) Ashkin-Teller model, in which two Ising spins, $\sigma^{\mathrm{A}}=\pm 1$ and $\sigma^{\mathrm{B}}=\pm 1$, exist at each lattice site. The four types of pair states, i.e., $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1),(-1,1),(-1,-1)$, and $(1,-1)$, correspond to the four states $s=0$, $1$, $2$, and $3$ of the Potts model, respectively. The phase behavior at thermal equilibrium has been investigated in detail [21, 23, 24]. Under the conditions $J_{k,k}>J_{k,[k+1]}$ and $J_{k,k}>J_{k,[k+2]}$, three phases exist: a disordered phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle=0$, a Baxter (paramagnetic) ordered phase with $\langle\sigma^{\mathrm{A}}\rangle\neq 0$, $\langle\sigma^{\mathrm{B}}\rangle\neq 0$, and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$, and a partially ordered phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=0$ and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$. In the partially ordered phase, two diagonal states ($s=0$ and $2$ or $s=1$ and $3$) are spatially mixed in the view of the Potts model, which corresponds to the M2 phase for the factorized number of states ($4=2\times 2$) in the present study.
Figure 18: Four-state Potts model at $L=128$. (a–d) Snapshots. The light-yellow, light-green, medium-blue, and red sites (light to dark in grayscale) represent $s=0$, $1$, $2$, and $3$, respectively. (a) Nonspiral waves (W4) at $h=1.1$ and $J_{k,[k+2]}=0$. (b) Spiral waves (W4) at $h=1$ and $J_{k,[k+2]}=-2$. (c) Homogeneous cycling of two diagonal states (HC2) at $h=1$ and $J_{k,[k+2]}=1$. (d) Mixing of two diagonal states (M2) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC4, HC2, M2, W4, and WI, respectively.
Figure 19: Dependence on $J_{k,[k+2]}$ in the four-state Potts model at $h=1$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. (b) Binder cumulant $U_{1}$ at $L=32$, $64$, and $128$. The bidirectional arrows at the top represent the ranges of dynamic modes.
Here, under active flipping of $h>0$, the M2 phase persists when $J_{k,[k+2]}\gtrsim 1$ (see Fig. 18(d) and (e)). For intermediate values of $J_{k,[k+2]}$, homogeneous cycling (switching) of two diagonal phases (HC2) emerges (see Fig. 18(c) and (e)). In the $k$-state dominant phase, the nuclei of $[k+1]$ state are swallowed up by the $[k+2]$ domains. In the M2 mode, no switching between even- and odd-numbered states occur in the considered simulation periods. These behaviors are similar to those of the M3 and HC3 modes of the six-state models. The transition between the HC2 and M2 modes is continuous, and the transition point is detected based on $U_{1}$ (see Fig. 19(b)). Since $R_{1}=0.57\pm 0.01$ at the transition point, we consider the system is taken as being in the M2 mode when $R_{1}<0.57$.
For $J_{k,[k+2]}\leq 0$, the HC4 and W4 modes manifest under low and high $h$, respectively. The WI mode appears only at $J_{k,[k+2]}\simeq-1$, and the direct HC4–W4 transition occurs through the temporal coexistence at $J_{k,[k+2]}\simeq 0$ and $J_{k,[k+2]}\simeq-2$ (see Figs. 18(e) and 19(a)). The waves form spiral shapes at $J_{k,[k+2]}\simeq-2$, whereas waves move separately without centers at $J_{k,[k+2]}\simeq 0$ (see Fig. 18(a) and (b)). The HC4 and HC2 modes are distinguished by the lifetime ratio and the state-density distribution, similar to the HC6 and HC3 modes in the six-state models.
III.4 Eight-State Potts Model
The eight-state Potts model corresponds to the three-color Ashkin-Teller model [22]. We obtained five modes at $J_{k,[k+2]}=J_{k,[k+3]}=J_{k,[k+4]}$: HC8, WI, W8, W4, and M8 (see Fig. 20(c)). The factorized number of states ($8=4\times 2$) are stabilized in the wave mode (W4), in which four domain types ($s=0,2,4$, and $6$ or $s=1,3,5$, and $7$) spatially coexist. In the W4 (W8) modes, the boundaries of the $k$ and $[k+2]$ ($[k+1]$) domains move in the direction from the $[k+2]$ ($[k+1]$) domain to the $k$ domains (see Fig. 20(a) and (b) and Movies S5 and S6).
The W8 and M8 modes can be distinguished by the contact probabilities $p_{\mathrm{contact}}$ or the mean cluster size $n_{\mathrm{c1}}$ (see Fig. 21(b) and (c)). Because $n_{\mathrm{c1}}$ exhibits a clear peak (or shoulder) at $J_{k,[k+2]}=0.4$, we conclude that the transition occurs at this peak point.
Figure 20: Eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $L=128$. (a) Snapshot of waves of eight states (W8) at $h=1.3$ and $J_{k,[k+2]}=0$. (b) Snapshot of waves of four states (W4) at $h=1$ and $J_{k,[k+2]}=0.6$. The white, yellow, green, light-blue, dark-blue, magenta, red, and black sites (light to dark in grayscale) represent $s=0$, $1$, $2$, $3$, $4$, $5$, $6$, and $7$, respectively. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC8, W4, M8, W8, and WI, respectively.
Figure 21: Dependence on $J_{k,[k+2]}$ in the eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $h=1.5$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, $k$$[k+3]$, and $k$$[k+4]$ pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=64$.
IV Relation with Other Models
The Avni model [38, 39] is an active four-state vector Potts model. They added a nonreciprocal interaction to the Ashkin-Teller model with two Ising spins ($\sigma^{\mathrm{A}}$and $\sigma^{\mathrm{B}}$). The energy of spin $\sigma^{\alpha}$ in the $i$-th site is given by
$E_{i}^{\alpha}=-J\sum_{\langle ij\rangle}\sigma_{i}^{\alpha}\sigma_{j}^{\alpha}-K\epsilon_{\alpha\beta}\sigma_{i}^{\alpha}\sigma_{i}^{\beta},$ (6)
where $J,K>0$ and $\epsilon_{\alpha\beta}$ is the Levi-Civita symbol. The energy of the first term decreases when the spins of the same type in the neighboring sites are in paramagnetic order. However, the second term gives a nonreciprocal intra-site interaction, such that the spin A is preferentially antimagnetic to the spin B at the same site ($\sigma_{i}^{A}\sigma_{i}^{B}<0$), whereas the spin B is preferentially paramagnetic ($\sigma_{i}^{A}\sigma_{i}^{B}>0$). Thus, the following cyclic change is induced: $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1)\to(-1,1)\to(-1,-1)\to(1,-1)\to(1,1)$. When the Avni model is mapped to the present active Potts model, $h=2K$, $J_{k,k}=2J$, $J_{k,[k+1]}=0$, and $J_{k,[k+2]}=-2J$ are obtained. The contact energy is expressed as $J_{k,[k+j]}=2J\cos(\pi j/2)$, such that these interactions correspond to those of the four-state vector Potts model (also called clock model) [16]. In the phase diagram of Fig. 18(e), the data obtained at $J_{k,[k+2]}=-2$ correspond to those of the Avni model at $J=1$. At low and high $h$, the HC4 and W4 modes occur, respectively, and the two modes temporally coexist around the transition point in the long-term dynamics.
Lattice Lotka–Volterra models, based on predator–prey interactions, have been widely used to study spatiotemporal patterns [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Cyclic changes in species (states) are considered similar to those of the present models. The Lotka–Volterra models of four or more species can exhibit short-term dynamics similar to those given by the present model, such as the long-lived spatial coexistence of two phases and domains of mixed species [41, 46, 47, 48, 49, 50, 51, 52, 53, 54]. However, the long-term dynamics completely differ. In the long-term limit, the Lotka–Volterra models exhibit only a uniform phase occupied by a single species through an absorbing transition. Extinct species never reappear, because species multiply through self-reproduction. Technically, this is proceeded by pair updates, such as $(s_{i},s_{j})=(0,1)\to(1,1)$, in which the $s_{i}$ species becomes the $s_{j}$ species, indicating predation and self-reproduction. These updates involve no backward processes; therefore, a detailed balance cannot be obtained. We compared the single update with this pair update using a three-state Potts model and demonstrated that the pair update generates fewer spirals at high flipping energy, but all sites are occupied by a single state in the long-term limit under all conditions (including $h=J_{ss^{\prime}}=0$[34]. Therefore, an update scheme including backward processes is crucial to obtain long-term spatiotemporal dynamics.
V Summary
We studied the nonequilibrium dynamics of the $q$-state active Potts models with $q=4$, $5$, $6$, and $8$ under cyclically symmetric conditions in the long-term limit. Three mode types and their combinations were identified: mixed phases, wave propagation modes, and cycling of homogeneous phases. In addition to these modes comprising $q$ states, modes comprising a factorized number of states emerge for $q=4$, $6$, and $8$: a mixed phase and homogeneous switching of two diagonal states for $q=4$, a wave mode of four states for $q=8$, and several modes for $q=6$. In the six-state model, two types of spiral waves with factorized symmetry were found: spiral waves of three states (three odd numbers, $s=1,3$, and $5$, or three even numbers, $s=0,2$, and $4$), with the other three states present as small clusters at the domain boundaries; and spiral waves of three types of mixed domains, in which two diagonal states ($s=k$ and $[k+3]$) are mixed within a domain. In the former and latter spirals, one flip cycle in the spiral domains requires one cycle and half cycle of states, respectively. Moreover, homogeneous cycling modes of the three odd- or even-numbered states and three mixed phases were found. Therefore, the factorization of symmetry can generate rich dynamic modes under nonequilibrium conditions, including the combinations of static (mixing) and dynamic (wave) modes.
We calculated the scaling exponents of the second-order transitions between the W3 and M6 modes and between the HC3 and M3 modes. The exponents of the W3–M6 transition are modified from the equilibrium values, whereas no change is detected for the HC3–M3 transition. The transitions to wave modes may generally exhibit greater changes in the exponents than those between static modes. In contrast, first-order transitions have not been found yet at $N\to\infty$ in the present models. Temporal coexistences of the HC and wave modes occur, but the transition points shift to $h=0$ with increasing system size. Since the forward and backward transitions typically take different pathways, they can exhibit different $N$ dependencies. Clarifying the condition for first-order transitions at $N\to\infty$ is an interesting problem for further studies.
In this study, we have kept the cyclic symmetry and fixed the temperature and contact energies between the same states and between neighboring states. In our previous studies, we reported non-cyclic homogeneous phases and the coexistence of two diagonal phases under asymmetric conditions (for three- and four-state Potts models) [35, 36]. For $q\geq 5$, the imposition of asymmetric conditions will likely modify the dynamic modes in a similar manner; however, symmetry factorization may additionally induce different types of dynamics. Even restricting to the symmetric conditions, we have not explored the entire parameter space. Under equilibrium, several antimagnetic phases have been reported for the Ashkin-Teller models [21, 22, 23, 24]. Thus, antimagnetic and other mode types likely exist in the remaining parameter space in active Potts models. Further, the modes and transitions can be changed in three (or higher)-dimensional space. Overall, the active Potts models provide a simple but rich system for exploring various types of nonequilibrium dynamics under thermal fluctuations.
Acknowledgements.
We thank Jun Takahashi for fruitful discussion. The simulations were partially carried out at ISSP Supercomputer Center, University of Tokyo (ISSPkyodo-SC-2025-Ca-0049). This work was supported by JSPS KAKENHI Grant Number JP24K06973.